OpenStudy (anonymous):

f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (potatoes.ramu):

I got 5 + f(4)

OpenStudy (anonymous):

How'd you get your answer?

OpenStudy (potatoes.ramu):

Substitute n=5 in the equation f(n)=f(1) +f(2)+f(5-1)

OpenStudy (potatoes.ramu):

We already know what f(1) and f(2) are.. Is it right?

OpenStudy (anonymous):

Ohhh ok! I get it now! Thank you so much! :)

OpenStudy (potatoes.ramu):

No problem!! :) Best of luck!!

hartnn (hartnn):

thats not correct, we can get a numerical answer for this

hartnn (hartnn):

start by finding f(3)

OpenStudy (anonymous):

How?

OpenStudy (potatoes.ramu):

Oh yeah! That's true!!

OpenStudy (potatoes.ramu):

Substitute f(3) in the equation I showed you. where n=3.

hartnn (hartnn):

to get f(3), just replace every 'n' by 3 in f(n) = f(1) + f(2) + f(n - 1)

OpenStudy (anonymous):

Ohhhh.

hartnn (hartnn):

so what u get f(3) as ?

OpenStudy (anonymous):

I got 5.. I don't think it's right...

hartnn (hartnn):

how did u get 5 ?

hartnn (hartnn):

f(n) = f(1) + f(2) + f(n - 1) f(3) = f(1)+f(2)+f(2) you know all values on right

OpenStudy (anonymous):

f(3)= 2+3+3=8 and use this to find f(4) the same way as hartnn mentioned f|(4)=f(1)+f(2)+f(3) and finally use it again to find f(5)